The Earth exerts a gravitational pull of 2*10^20 N on the Moon;the sun exerts a gravitational pull of 4.3*10^20 N on the Moon.What is the net force on Moon ; when the angular separation between the Earth and the Sun is 90 as seen from the Moon?
1
Expert's answer
2019-09-18T09:49:30-0400

Solution:

Net force on Moon:

"F=\sqrt{F^2_1+F^2_2}=\sqrt{(2\sdot{}10^{20})^2+(4.3\cdot{}10^{20})^2}="

"=\sqrt{(2)^2+(4.3)^2}\cdot10^{20}=\sqrt{22.49}\cdot10^{20}="

"=4.7\cdot10^{20}" N.

Answer:

Net force on Moon is "4.7\cdot10^{20}" N.